Max is 24 years older than his son, Liam. In two years, Liam will be half as old as Max. How old is Liam now?
Answer by algebrahouse.com(1659)

"Max is 24 years older than his son, Liam. In two years, Liam will be half as old as Max. How old is Liam now?"

x = Liam's age now
x + 24 = Max's age now

x + 2 = Liam's age in 2 years
x + 26 = Max's age in 2 years

x + 2 = 1/2(x + 26) {in two years, Liam will be half as old as Max}
x + 2 = 1/2x + 13 {used distributive property}
x = 1/2x + 11 {subtracted 2 from both sides}
1/2x = 13 {subtracted 1/2x from both sides}
x = 22 {divided both sides by 1/2 or multiplied by reciprocal, 2}

Liam is 22 now.
www.algebrahouse.com